OpenStudy (anonymous):

A man of height 2.1 meters walk away from a 5-meter lamppost at a speed of 1.5 m/s. Find the rate at which his shadow is increasing in length

OpenStudy (anonymous):



OpenStudy (anonymous):

By similar triangle with y is the shadow length: y/ 2.1 = ( x + y ) / 5

OpenStudy (anonymous):

@heyheyhey111 Questions?

OpenStudy (anonymous):

uhm let me work it out thanks I will ask again if I have questions

OpenStudy (anonymous):

Okay uhm, how do I solve for just x, or just y? I tried to solve it out and I got y =2.1x / 2.9

OpenStudy (anonymous):

actually would x be equal to the 1.5?

OpenStudy (anonymous):

Yes haha I got it right, kind of talking to myself but you helped me with the triangles thanks!

OpenStudy (anonymous):

y =2.1x / 2.9 "Find the rate at which his shadow is increasing in length" at x' = 1.5 m/s. -> y' = .... Have you got to y' ?